The time taken to complete a particular type of job is distributed approximately normally with mean 1.8 hours and standard deviation 0.1 hours. (a) If normal-time work finishes at 6.00 p.m. and a job is started at 4.00 p.m., what is the probability that the job will need overtime payments? (b) What estimated completion time (to the nearest minute) should be set so that there is a 90% chance that the job is completed on time?


1
Expert's answer
2022-05-20T10:50:48-0400

Given "\mu=1.8, \sigma=0.1"

"P(X>2)=1-P(Z\leq\dfrac{2-1.8}{0.1})"




"=1-P(Z\leq2)\approx0.02275"




"P(X<x)=P(Z<\dfrac{x-1.8}{0.1})=0.9""\dfrac{x-1.8}{0.1}\approx1.6449""x=1.8+0.16449"

"x=1.96449\ h"


"x=1hr\ 58\ min"

The estimated completion time is 5.58 pm.